OpenStudy (anonymous):

(3 x 10^5) x (2 x 10^6)

geerky42 (geerky42):

Remember, \(\large a^b a^c = a^{b+c}\).

OpenStudy (anonymous):

Multiply the 2 and 3 and the powers of 10 to get 6 x 10^11

geerky42 (geerky42):

For explain, \(\large 5^2 5^3 = 5^{2+3} = 5^5\)

OpenStudy (anonymous):

Ahh yes, I remember now.